OpenStudy (anonymous):

What does x equal? 3/4 x + 5/8=4x

OpenStudy (anonymous):

extend 4x on the RHS into a fraction with /4

OpenStudy (anonymous):

Whats RHS?

OpenStudy (anonymous):

Right hand side :)

OpenStudy (anonymous):

Okay :))) ty

OpenStudy (anonymous):

4x can be written as this right?\[\frac{ 4x }{ 1 }\]

OpenStudy (anonymous):

yeah im writing that down now.. and you multiply it with the other fractions? correct?

OpenStudy (anonymous):

on the left side, there's 3/4 x so, to add or subtract the two x terms, the terms must have the same denominator

OpenStudy (anonymous):

\[\frac{ 3 }{ 4 }x + \frac{4}{1} x \neq \]

OpenStudy (anonymous):

So it'll be set up as \[\frac{ 3 }{ 4 }x + \frac{ 5 }{ 8 }= \frac{ 4x }{ 1 }\]

OpenStudy (anonymous):

Oh never mind

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it will?

OpenStudy (anonymous):

right, and you can extend the 4x term so that it will have a denominator of 4

OpenStudy (anonymous):

for following operations

OpenStudy (anonymous):

we want all x terms to have the same denominator

OpenStudy (anonymous):

alright im trying to do that now. sorry its a bit confusing for me.. i hate fractions

OpenStudy (anonymous):

Im really confused right now omg..

OpenStudy (anonymous):

fractions: \[\frac{ 1 }{ 1 } = \frac{ 4 }{ 4 } = \frac{ 9 }{ 9 }\]right?

OpenStudy (anonymous):

Yessss

OpenStudy (anonymous):

and also \[\frac{ 1 }{ 2 } = \frac{ 2 }{ 4 } = \frac{ 4.5 }{ 9 }\]

OpenStudy (anonymous):

mhmmmmmm

OpenStudy (anonymous):

so you're saying i have to make the fractions like that pretty much but with the same denominator, right?

OpenStudy (anonymous):

we have a fraction on the left, no fraction on the right. so the non-fractional variable must be made a fractional variable to add or subtract them. also, the one on the left has a 4 in the denominator. because we can only add/subtract fractions that have THE SAME denominator, the x variable on the right should get a 4 denominator.

OpenStudy (anonymous):

it should get a 4 as denominator, but the value must remain the same

OpenStudy (anonymous):

the value of the term may not change, only its form

OpenStudy (anonymous):

\[\frac{ 3 }{ 4 }x +\frac{ 7 }{ 4 }x\] like that?

OpenStudy (anonymous):

where do you get 7/4 x ?

OpenStudy (anonymous):

You said to change the fraction not the value but the term so i put 4/1 as 7/4 ... nevermind its fine i dont wanna waste your time im dumb sorry :/

OpenStudy (anonymous):

well looks like it's the right idea :) however 4/1 = 4 and 7/4 = 1.75 so....they're not the same at all

OpenStudy (anonymous):

to keep the same value you can take the original one: 4/1 and them multiply both denominator and numerator with the same number. if you do that, the value remains the same.

OpenStudy (anonymous):

I made \[\frac{ 1 }{ 4}\] into \[\frac{ 8 }{ 4}\]

OpenStudy (anonymous):

*4/1

OpenStudy (anonymous):

1/4 = 0.25 8/4 = 2.00

OpenStudy (anonymous):

sorry i meant 4/1

OpenStudy (anonymous):

ok, if you make 4/1 into 8/4 then it is half the value as before, because you multiply the above by 2 and the below by 4 so since your multiplication has a difference of 2/4 or 1/2 the value is half. 4/1=4 8/4=2

OpenStudy (anonymous):

Okay so i multiplied 4 x 4 and 1 x 4 and got 8/4 but it says 8/4 = 2 and 4/1 = 4

OpenStudy (anonymous):

yeah because 4x4 = 16 :)

OpenStudy (anonymous):

OH MY GOD 0EIUWHFR

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

HAHAHAH SORRY IM BRAIN DEAD IM SLEEPY IM DUMB SORRY

OpenStudy (anonymous):

I WAS THINKING 4+4

OpenStudy (anonymous):

okay so 4/1 = 16/4

OpenStudy (anonymous):

no problem. and yes: this is what we have so far\[\frac{ 3 }{ 4 }x + \frac{ 5 }{ 8 }= \frac{ 16 }{ 4 }x\]

OpenStudy (anonymous):

YAY I DID SOMETHINGGG :)))))) and dont i have to change the 5/8 or no?

OpenStudy (anonymous):

we can leave it just now... since we're not allowed to combine x and non-x terms anywy

OpenStudy (anonymous):

Okay so, its 3/4 x + 5/8 + 16/4 x and dont i have to subtract 5/8 from 16/4?

OpenStudy (anonymous):

you can only subtract x terms from each other and non-x terms from each other we must subtract 3/4 x from 16/4 x

OpenStudy (anonymous):

because they're the only terms that have the same "type" (x)

OpenStudy (anonymous):

I got -13/4x

OpenStudy (anonymous):

you subtracted 16/4x from 3/4 x ; )

OpenStudy (anonymous):

MY MOM TOLD ME TO PUT IT THAT WAY OMG LOL

OpenStudy (anonymous):

ok maybe I'm wrong on this one

OpenStudy (anonymous):

add 2 to 3 3+2 subtract 3 from 5 5-3

OpenStudy (anonymous):

its fine, i do lots of stuff wrong

OpenStudy (anonymous):

anyway if we want a positive number we would do 16/4x - 3/4x = 13/4x

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

and if we do this on both sides then we got rid of the 3/4x on the left: \[\frac{ 5 }{ 8 }= \frac{ 13 }{ 4 }x\]

OpenStudy (anonymous):

because\[\frac{ 3 }{ 4 }-\frac{ 3 }{ 4 }=0\]

OpenStudy (anonymous):

and we divide it right?

OpenStudy (anonymous):

we can't divide 13 by 4, but we want to solve for x which means we want x alone as one term in the end, x=... \[\frac{ 5 }{ 8 }= \frac{ 13 }{ 4 }x\]

OpenStudy (anonymous):

we can multiply both sides by 4, to get rid of the fraction in x

OpenStudy (anonymous):

\[\frac{ 13 }{ 4 }x * 4\]is the same as this one:\[13x\]

OpenStudy (anonymous):

so i multiply 4 times 13 and 4 and use that fraction?

OpenStudy (anonymous):

which is 52/16

OpenStudy (anonymous):

you did a correct transformation and conserved the value. My suggestion is different and aims to simplify the x term and actually changing its value. taking the x term 4 times will take away the 4 in the denominator, and change the value. this is allowed IF we also change the value on the left side the same way. so if we multiply the left side by 4, then we are allowed to take the x term 4 times and taking the x term 4 times simply removes the denominator. \[\frac{ 5 }{ 8 }= \frac{ 13 }{ 4 }x |\times 4\]\[\frac{ 20 }{ 8 }= 13x\]

OpenStudy (anonymous):

I have to divide next, correct?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

dividing a fraction by a number is the same thing as multiplying its denominator with that number example: to divide 1 by 2, we can also write: \[\frac{ 1 }{ 1 \times 2 }=\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

the same law works when you divide 20/8 by 13

OpenStudy (anonymous):

i got 20/104

OpenStudy (anonymous):

yep, I also got that :) x = 20/104 I'm not sure if you can simplify the fraction. probably not

OpenStudy (anonymous):

ok you can already divide by 2

OpenStudy (anonymous):

I DID IT OMG TYSM YOU ARE JESUS

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

TYYYYYYSSSSSMMM IT MEANS A LOT

OpenStudy (anonymous):

x = 5/26

OpenStudy (anonymous):

YEAH I GOT THAT HA

OpenStudy (anonymous):

ok have a nice day ;)

OpenStudy (anonymous):

YOU TOOOOO

OpenStudy (anonymous):

thanks